OpenStudy (dumbsearch2):

The sector shows the area of a lawn that will be watered by a sprinkler. What is the area, rounded by the nearest tenth? Use 3.14 for pi.

OpenStudy (dumbsearch2):

The sector shows the area of a lawn that will be watered by a sprinkler. What is the area, rounded by the nearest tenth? Use 3.14 for pi.

OpenStudy (kohai):

I learned this a few years ago, let me do some sniffing

jimthompson5910 (jim_thompson5910):

Which degree measure corresponds to a full rotation?

OpenStudy (kohai):

Or jim can help you out ;)

OpenStudy (dumbsearch2):

Wait, 360? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes 360 degrees

jimthompson5910 (jim_thompson5910):

what fraction of 360 is 30?

OpenStudy (dumbsearch2):

12

jimthompson5910 (jim_thompson5910):

more like 1/12 actually

OpenStudy (dumbsearch2):

But how could I calculate the area from here?

jimthompson5910 (jim_thompson5910):

since 30/360 = 1/12

jimthompson5910 (jim_thompson5910):

so 1/12 of the circumference is equal to 10 ft, which is the arc length

jimthompson5910 (jim_thompson5910):

what is the circumference of this whole circle?

OpenStudy (dumbsearch2):

@jim_thompson5910 isn't the circumference 10x12?

jimthompson5910 (jim_thompson5910):

120 yep

jimthompson5910 (jim_thompson5910):

now you know the circumference

jimthompson5910 (jim_thompson5910):

how can you use this to find the radius of the circle?

OpenStudy (dumbsearch2):

Yes but the area formula does not involve the circumference, so how do we calculate the radius from there?

jimthompson5910 (jim_thompson5910):

what is the circumference formula?

OpenStudy (dumbsearch2):

pi*d

jimthompson5910 (jim_thompson5910):

and d = 2r

OpenStudy (dumbsearch2):

yes

jimthompson5910 (jim_thompson5910):

so C = pi*2r or C = 2*pi*r

jimthompson5910 (jim_thompson5910):

you can use C = 120 to find r

OpenStudy (dumbsearch2):

but I don't know how to calculate the radius

jimthompson5910 (jim_thompson5910):

plug in C = 120 and solve for r

OpenStudy (dumbsearch2):

I finished the process and I found the answer!! 95 Thank you so much for guiding me through the steps.